# Argumentative Large Language Models for Explainable and Contestable Claim Verification

### v2

@@ 1 Introduction
Throughout the paper, we focus on claim verification as the decision-making task of interest. This setting lends itself well to ArgLLMs as claims are often under-determined, so they may not have straightforward truth values. By intrinsically considering both arguments in favour of and in conflict with the truthfulness of claims, ArgLLMs are able to ascertain the best answer given the available evidence. For simplicity, and without loss of generality, [^1] ArgLLMs can be easily extended to the case where there are multiple possible options by generating a set of candidate answers and applying our method to each of these. we focus on the binary setting, rather than general question-answering as in Wei et al. (2022) and Yao et al. (2023).
@@ 3 Preliminaries
A QBAF Baroni et al. (2019) is a quadruple $\langle\mathcal{A},\mathcal{R}^{-},\mathcal{R}^{+},\tau\rangle$ comprising: a set of arguments $\mathcal{A}$; binary, directed relations of attack $\mathcal{R}^{-}\!\subseteq\!\mathcal{A}\!\times\!\mathcal{A}$ and support $\mathcal{R}^{+}\!\subseteq\!\mathcal{A}\!\times\!\mathcal{A}$, where $\mathcal{R}^{-}\!\cap\!\mathcal{R}^{+}=\emptyset$; and a total function $\tau\!:\!\mathcal{A}\!\rightarrow\![0,1]$, where for any $\alpha\in\mathcal{A}$, $\tau(\alpha)$ is the base score of $\alpha$ [^2] The codomain of $\tau$ is defined more generally by Baroni et al. (2019), but we use here its most common form. For any argument $\alpha\in\mathcal{A}$, we use $\mathcal{R}^{-}(\alpha)\!=\!\{\beta\!\in\!\mathcal{A}|(\beta,\alpha)\!\in\!\mathcal{R}^{-}\}$ to refer to the attackers of $\alpha$ and $\mathcal{R}^{+}(\alpha)\!=\!\{\beta\!\in\!\mathcal{A}|(\beta,\alpha)\!\in\!\mathcal{R}^{+}\}$ to refer to the supporters of $\alpha$. Arguments in QBAFs may be evaluated by a gradual semantics Baroni et al. (2019), i.e. a total function $\sigma:\mathcal{A}\rightarrow[0,1]$ which, for any $\alpha\in\mathcal{A}$, assigns a strength $\sigma(\alpha)$ to $\alpha$ [^3] As with $\tau$, we use the most common codomain for $\sigma$. While $\tau(\alpha)$ can be seen as the intrinsic strength for $\alpha$, the strength $\sigma(\alpha)$ can be seen as ‘dialectical’, following the debate captured by $\mathcal{R}^{-}$ and $\mathcal{R}^{+}$. If we ignore $\tau$ in a QBAF, we obtain a bipolar argumentation framework (BAF) Cayrol and Lagasquie-Schiex (2005).
@@ 5 Performance Evaluation
In this section, we describe the experiments evaluating the reasoning performance of ArgLLMs on our selected task of claim verification. [^4] All our experiments are executed with two RTX 4090 24GB GPUs on an Intel(R) Xeon(R) w5-2455X. We derive claims from existing QA datasets (for an example, see Figure 1). Also, we experiment with a range of parameters, $\theta$, in the argument generation component and with several choices for $E$. We report results for DF-QuAD as $\sigma$ in the $\Sigma$ component, but include results for another choice of gradual semantics (quadratic energy Potyka (2018)) in the Supplementary Material in the extended version (referred to in the remainder in short as SM).
@@ 5.1 Experimental Set-up
We use seven main models: Mistral (Mistral-7B-Instruct-v0.2) Jiang et al. (2023), Mixtral (Mixtral-8x7B-Instruct-v0.1) Jiang et al. (2024), Gemma (gemma-7b-it) Mesnard et al. (2024), Gemma 2 (gemma-2-9b-it) Riviere et al. (2024), Llama 3 (Meta-Llama-3-8B-Instruct) Dubey et al. (2024), GPT-3.5-turbo (GPT-3.5-turbo-0125) Brown et al. (2020) and GPT-4o mini (gpt-4o-mini) OpenAI (2024). We chose Mistral, Mixtral, Gemma and Llama 3 as they were among the most known and best-performing open-source [^5] We consider a broad notion of the term “open-source”, not necessarily implying the use of OSI-approved licenses. models of reasonable size at the time of our evaluation. In order to reduce the computational costs of running the open-source models, we quantise them to 4 bits Dettmers et al. (2023) for both the baselines and our method. As representatives of models with proprietary weights, we chose GPT-3.5-turbo and GPT-4o mini, since they had the best performance/cost trade-off. For all the models, we use parameters: temperature 0.7, max new tokens for arguments 128, max new tokens for baselines 768, top-p 0.95 and repetition penalty 1.0.
@@ 6 Contestability
This property formalises that increasing the base score of pro/con arguments in a QBAF for an argument will not decrease/increase (respectively) the argument’s final strength [^6] Pro/con arguments no longer affect $\sigma_{\mathcal{Q}}(\alpha^{*})$ when it is 1 or 0..

### v3

@@ 1 Introduction
Throughout the paper, we focus on claim verification as the decision-making task of interest. This setting lends itself well to ArgLLMs as claims are often under-determined, so they may not have straightforward truth values. By intrinsically considering both arguments in favour of and in conflict with the truthfulness of claims, ArgLLMs are able to ascertain the best answer given the available evidence. For simplicity, and without loss of generality,ArgLLMs can be easily extended to the case where there are multiple possible options by generating a set of candidate answers and applying our method to each of these. we focus on the binary setting, rather than general question-answering as in Wei et al. (2022) and Yao et al. (2023).
@@ 3 Preliminaries
A QBAF Baroni et al. (2019) is a quadruple $\langle\mathcal{A},\mathcal{R}^{-},\mathcal{R}^{+},\tau\rangle$ comprising: a set of arguments $\mathcal{A}$; binary, directed relations of attack $\mathcal{R}^{-}\!\subseteq\!\mathcal{A}\!\times\!\mathcal{A}$ and support $\mathcal{R}^{+}\!\subseteq\!\mathcal{A}\!\times\!\mathcal{A}$, where $\mathcal{R}^{-}\!\cap\!\mathcal{R}^{+}=\emptyset$; and a total function $\tau\!:\!\mathcal{A}\!\rightarrow\![0,1]$, where for any $\alpha\in\mathcal{A}$, $\tau(\alpha)$ is the base score of $\alpha$.The codomain of $\tau$ is defined more generally by Baroni et al. (2019), but we use here its most common form. For any argument $\alpha\in\mathcal{A}$, we use $\mathcal{R}^{-}(\alpha)\!=\!\{\beta\!\in\!\mathcal{A}|(\beta,\alpha)\!\in\!\mathcal{R}^{-}\}$ to refer to the attackers of $\alpha$ and $\mathcal{R}^{+}(\alpha)\!=\!\{\beta\!\in\!\mathcal{A}|(\beta,\alpha)\!\in\!\mathcal{R}^{+}\}$ to refer to the supporters of $\alpha$. Arguments in QBAFs may be evaluated by a gradual semantics Baroni et al. (2019), i.e. a total function $\sigma:\mathcal{A}\rightarrow[0,1]$ which, for any $\alpha\in\mathcal{A}$, assigns a strength $\sigma(\alpha)$ to $\alpha$.As with $\tau$, we use the most common codomain for $\sigma$. While $\tau(\alpha)$ can be seen as the intrinsic strength for $\alpha$, the strength $\sigma(\alpha)$ can be seen as ‘dialectical’, following the debate captured by $\mathcal{R}^{-}$ and $\mathcal{R}^{+}$. If we ignore $\tau$ in a QBAF, we obtain a bipolar argumentation framework (BAF) Cayrol and Lagasquie-Schiex (2005).
@@ 5 Performance Evaluation
In this section, we describe the experiments evaluating the reasoning performance of ArgLLMs on our selected task of claim verification.All our experiments are executed with two RTX 4090 24GB GPUs on an Intel(R) Xeon(R) w5-2455X. We derive claims from existing QA datasets (for an example, see Figure 1). Also, we experiment with a range of parameters, $\theta$, in the argument generation component and with several choices for $E$. We report results for DF-QuAD as $\sigma$ in the $\Sigma$ component, but include results for another choice of gradual semantics (quadratic energy Potyka (2018)) in the Supplementary Material in the extended version (referred to in the remainder in short as SM).
@@ 5.1 Experimental Set-up
We use seven main models: Mistral (Mistral-7B-Instruct-v0.2) Jiang et al. (2023), Mixtral (Mixtral-8x7B-Instruct-v0.1) Jiang et al. (2024), Gemma (gemma-7b-it) Mesnard et al. (2024), Gemma 2 (gemma-2-9b-it) Riviere et al. (2024), Llama 3 (Meta-Llama-3-8B-Instruct) Dubey et al. (2024), GPT-3.5-turbo (GPT-3.5-turbo-0125) Brown et al. (2020) and GPT-4o mini (gpt-4o-mini) OpenAI (2024). We chose Mistral, Mixtral, Gemma and Llama 3 as they were among the most known and best-performing open-sourceWe consider a broad notion of the term “open-source”, not necessarily implying the use of OSI-approved licenses. models of reasonable size at the time of our evaluation. In order to reduce the computational costs of running the open-source models, we quantise them to 4 bits Dettmers et al. (2023) for both the baselines and our method. As representatives of models with proprietary weights, we chose GPT-3.5-turbo and GPT-4o mini, since they had the best performance/cost trade-off. For all the models, we use parameters: temperature 0.7, max new tokens for arguments 128, max new tokens for baselines 768, top-p 0.95 and repetition penalty 1.0.
@@ 6 Contestability
This property formalises that increasing the base score of pro/con arguments in a QBAF for an argument will not decrease/increase (respectively) the argument’s final strengthPro/con arguments no longer affect $\sigma_{\mathcal{Q}}(\alpha^{*})$ when it is 1 or 0..